# Reinforcement Learning in Non-Markov Market-Making

### v1

@@ 1 Introduction
The last 20 years has seen the rising dominance of algorithmic and High-Frequency Trading (HFT) in some of the most liquid financial markets, as this method has become the most common way to complete trade transactions. Transactions in some of the these highly liquid markets, such as equities, futures, or currencies, in particular, occur through the so-called limit order book (LOB) mechanism, which connects the buyers and sellers for the tradeable assets in these financial markets. See figure 1, below, where we show a visual description of the LOB in the E-mini S&P 500 contract (ES) on April 24th, 2024, which shows a snapshot at a random time during the most active US stock market trading hours (9:30 EST - 16:00 EST). Here one can see the bid and ask prices displayed in the LOB, as well as their respective queue sizes. For a very informed and broad survey on how LOBs are modelled, see Gould et al., (2013), where they highlight many of the major findings from the theoretical and empirical literature on LOBs, as well as addressing many of their limitations, with many of the details in the literature still inadequately formulated. The algorithms used to trade in these markets can place a whole variety of trade order types, where we will specifically focus on strategies that solely implement market and limit orders for simplicity purposes. Extensions can quite often easily be made for the wide array of different trade order types.
@@
The modelling of stochastic processes in mathematical finance has been a major topic for many years, where now more attention is being placed on modelling the variables in algorithmic and HFT markets. Stochastic Optimal Control (SOC) theory has often been the main method for tackling trading problems in this realm, where Cartea et al., (2015) formulates a large number of examples deemed relevant for different types of trading needs. Their models often begin with a simple price process following a general arithmetic Brownian motion model. These price processes, however, have been proven to be less well-suited within the algorithmic and HFT setting, as many studies have shown that LOB dynamics often follow non-Markovian properties. One obvious oversight that the literature has found by studying LOB data is that it often experiences jumps i.e., points of discontinuity. The general conclusion from many studies is that LOB models that can portray a dependency in past trade transactions is superior to models with the assumption of an infinitesimal tick size seen in the arithmetic Brownian motion model, as can be seen from the empirical results in He and Swishchuk, (2019), Swishchuk et al., (2019) and Swishchuk and Huffman, (2020). Many more studies have provided similar evidence, where some in-depth examples can be found in Cartea et al., 2018b , Cartea et al., 2018a , Mäkinen et al., (2019), Sjogren and DeLise, (2021), Gašperov and Kostanjčar, (2022) and many more.
As well as improving the price process modelling of LOBs, recent major advancements in Reinforcement Learning (RL), particularly in Deep RL, have enabled the formulation of a more state-of-the-art framework for solving problems in algorithmic and HFT. The most commonly used framework for studying these types of problems has normally been the AS model from Avellaneda and Stoikov, (2008), with numerous examples for various different types of trading problems such as liquidation, acquisition, MM, pairs-trading and statistical arbitrage given in Cartea et al., (2015) under the SOC setting. These methods, however, are not very robust and an alternative approach that has recently increased in popularity would be to find near-optimal controls under a deep RL framework, which is essentially an approximate solution deviating from the optimal solution by a small amount.
@@ 2 Deep Reinforcement Learning
SOC has often been the standard framework to solve algorithmic and HFT trading problems, with some important works including Bertsimas and Lo, (1998) in dealing with execution costs, Bouchard et al., (2011) where a general impulse control approach is applied to algorithmic trading problems, Fodra and Pham, (2015) applies optimal control to a to a HFT problem under a Markov Renewal approach, Cartea et al., (2015) gives a whole range of algorithmic and HFT problems such as acquisition, liquidation, MM, volume imbalance, statistical arbritrage, pairs trading, and in Guéant, (2017) a thorough theoretical overview is given for the optimal MM problem. In more recent times, the focus has shifted from SOC to deep RL for solving problems in algorithmic trading and HFT. One can find some recent general trading examples with applications in Li et al., (2019), Zhang et al., (2019), Théate and Ernst, (2021), and Cartea et al., (2021). This shift is mainly due to the inherent advantages in the RL space and we will now discuss three main reasons related to the kind of problems seen in algorithmic and HFT:
@@
    Model Uncertainty: SOC generally requires a well-defined model of the market dynamics, transition probabilities and reward structure. However, in reality, modelling financial markets is extremely complex due to the presence of many latent variables which could include sentiment, order flow information, liquidity, market microstructure noise, trading interactions and many more. RL does not require a well-defined model of these dynamics, and it can learn optimal strategies from interacting directly with the market. Thus, RL learns to continuously adapt to the markets highly unpredictable stochastic nature, which is essential for developing dynamic systems requiring flexible models.
@@ 4 Results
In order to perform our analysis we first trained our model on our simulated data, which was constructed based on the analysis in section 3. Please see below, in table 1, the parameter values used. The majority of these parameters were purely picked for illustrative purposes to assess how the deep RL algorithm could learn and perform, although some strategic selections were made. We kept the time step small as we would like to assess the strategy from a HFT standpoint. We set the maximum inventory parameter $q=5$ to avoid large positions significantly influencing the results. We picked $20\%$ for our non-adverse fill probability, which is in line with empirical results in Lalor and Swishchuk, 2024b for estimating this probability based on some of the most liquid futures contracts listed on the Chicago Mercantile Exchange. We set the spread between the bid and the ask to $\Delta=0.01$ which is quite common throughout the majority of the day in many liquid financial products and we set the running inventory penalty to $0.001$ which had quite a bit of emphasis on the test results as inventory is generally very low. For our in-sample results, we set the number of simulations to $10^{6}$ for the training set, which lead to us training 1000 episodes. For our out-of-sample results, we then tested our training results on 200 unseen episodes. Computationally, we performed our analysis on python and for the SAC analysis part we made use of the Stable Baselines3 python package, developed by Raffin et al., (2020), which was also used by Gašperov and Kostanjčar, (2022) in their optimal MM problem.
@@
To summarize the training results, we provide visuals in figures 3 and 4. In the training phase, the agent interacts with the trading environment by taking actions based on the policy. Following every action the MM agent takes, a reward is received after which they transition to the next state. The agents policy is then updated based on its reward function and its state transitions through the SAC algorithm, which we described in section 3. In figure 3, we show some results over the first training episode to portray how the optimal MM strategy evolved there. In the top subplot, we show a snapshot of the simulated bid and ask prices, along with the trade executions, then we show the simulated inventory in the second subplot, and in the last subplot we show the simulated cumulative reward. Unsurprisingly, this agent has a negative cumulative reward at the beginning of the training process as it has not been able to figure out the optimal actions to take yet. As we get through more episodes, this slowly improves. Then to portray the cumulative results over all 1000 episodes, see the histogram in figure 4, which shows the frequency of every terminal reward over all these episodes. It is quite apparent that the results are mostly positive, with some heavy left tail results with extreme negative rewards. This is likely influenced by a large number of adverse fills and through the non-adverse fill probabilities, as it created a more realistic scenario where these types of unfavorable events are more likely to occur.
@@
Next, we discuss the results from our test sample, $X_{test}$, which was conducted over 200 episodes. This step is crucial as it helps us determine whether our agent has learned to perform in this type of trading environment and whether a similar performance can be achieved on out-of-sample data. During the testing phase, the agent’s policy is fixed, which means it does not update its policy based on any rewards received. The environment we used under our training and testing sets are the same, which is based on our analysis in section 3, and this ensures unbiased results in machine learning terms. However, to assess the real-world deployment capabilities of any optimal MM, it would be prudent to also test the results of the training model in different types of market regimes, in particular regimes that were unseen in the training data, as this could significantly alter the out-of-sample results. Markets regularly change from one regime to another, which can often be measured through metrics such as the implied volatility. One should also test the robustness of any optimal MM strategy like this one, which can be done by changing the parameter values given in table 1.
@@
To visualize the testing results, we now provide similar visuals as we just did above for the training results, in figures 5 and 6. It is quiet obvious that the results have significantly improved, in particular regarding the left tail events present in the training phase, and the histogram shows that the cumulative reward was positive in each episode. In figure 5, we again show a snapshot of the simulated bid and ask prices in the top subplot, along with the trade executions, then we show the simulated inventory in the second subplot, and in the last subplot we show the simulated cumulative reward. One thing that is very apparent is that the deep RL algorithm has learned it is best to not let the inventory deviate from -1 or 1. Thus, keeping a small position is optimal in this scenario. This likely helped to reduce the left tail events as the variables such as adverse fills have less of an effect when they occur if the inventory is low, while this outcome was also likely influenced by the inventory penalty term, $\alpha$. In figure 6, we can see that the cumulative terminal reward after each episode is positive, thus under this specific framework the algorithm has learned to consistently generate positive rewards.
@@
Figure 3: Training results from the first training episode.
@@
Figure 4: Histogram of cumulative rewards over all 1000 training episodes.
@@
Figure 5: Testing results from the last testing episode.
@@
Figure 6: Histogram of cumulative rewards over all 200 testing episodes.
@@
As good as the above results may seem (in terms of the cumulative rewards), we would like to stress some of their limitations, as much recent work has shed some more light on the limitations of trading strategy back-test results under diffusion pricing models, in particular in a setting where many limit orders are placed, as in this MM strategy. We will now list some of the general limitations, where a great overview was given in Law and Viens, (2019). We did try to overcome some of these limitations by making adjustments to our optimal MM strategy, which we will also highlight again. These limitations and adjustments are as follows:
@@
  • In these types of frameworks, price and order arrivals are assumed to be independent. In order to tackle this problem, we made sure to include adverse fills in the simulation process as in Lalor and Swishchuk, 2024b . Excluding adverse fills has the ability to lead to “large phantom gains”, as described in Law and Viens, (2019), as MMs as liquidity providers are often on the wrong side of the trade. The empirical results in section 1.1 in Lalor and Swishchuk, 2024b also shows this. Thus, including adverse fills has the ability to bring the results a lot closer to reality but still with some limitations, as it doesn’t exactly mimic how trading in live markets actually occurs, which is purely discrete based on specific LOB events. An overview of these events, under a reduced-form LOB model, can be seen in table 2 in Law and Viens, (2019), which was also applied to the discrete non-diffusive optimal MM deep RL problem in Gašperov and Kostanjčar, (2022).
@@ 5 Conclusions and Future Recommendations
In terms of recommendations for future research, we would like to highlight that if one could employ a full LOB model into the trading environment, as briefly described in Law and Viens, (2019), this would be a big step forward in improving the applicability of these types of models. In Gašperov and Kostanjčar, (2022) an attempt is made to create a reduced-form LOB model, where they also cite Law and Viens, (2019), which deploys a discrete-time multivariate Hawkes process model based on eight of the twelve events in the reduced-form LOB system. Other, and probably more simple model recommendations that we would recommend includes developing models where the spread is not fixed, computing stochastic non-adverse fill probabilities and stochastic volatility coefficients related to all the jump-diffusion coefficients, which may be related to either the semi-Markov or Hawkes models. In terms of the different types of trading problems out there, one could also try to apply this deep RL setting to the trading problems related to Liquidation, Acquisition, Statistical Arbitrage, Volume Imbalance and Pairs Trading as examples, which are some of the heavily studied trading problems under the standard SOC framework in Cartea et al., (2015).

### v2

@@ 1 Introduction
The last 20 years has seen the rising dominance of algorithmic and High-Frequency Trading (HFT) in some of the most liquid financial markets, as this method has become the most common way to complete trade transactions. Transactions in some of the these highly liquid markets, such as equities, futures, or currencies, in particular, occur through the so-called limit order book (LOB) mechanism, which connects the buyers and sellers for the tradeable assets in these financial markets. See figure 1, below, where we show a visual description of the LOB in the E-mini S&P 500 contract (ES) on April 24th, 2024, which shows a snapshot at a random time during the most active US stock market trading hours (9:30 EST - 16:00 EST). Here one can see the bid and ask prices displayed in the LOB, as well as their respective queue sizes. For a very informed and broad survey on how LOBs are modelled, see Gould et al., (2013), and for a more recent survey paper on simulating LOBs see Jain et al., (2024), where they highlight many of the major findings from the theoretical and empirical literature on LOBs, as well as addressing many of their limitations, with many of the details in the literature still inadequately formulated. The algorithms used to trade in these markets can place a whole variety of trade order types, where we will specifically focus on strategies that solely implement market and limit orders for simplicity purposes. Extensions can quite often easily be made for the wide array of different trade order types.
@@
The modeling of stochastic processes in mathematical finance has been a major topic for many years, where now more attention is being placed on modelling the variables in algorithmic and HFT markets. Stochastic Optimal Control (SOC) theory has often been the main method for tackling trading problems in this realm, where Cartea et al., (2015) formulates a large number of examples deemed relevant for different types of trading needs. Their models often begin with a simple price process following a general arithmetic Brownian motion model. These price processes, however, have been proven to be less well-suited within the algorithmic and HFT setting, as many studies have shown that LOB dynamics often follow non-Markovian properties. One obvious oversight that the literature has found by studying LOB data is that it often experiences jumps i.e., points of discontinuity. The general conclusion from many studies is that LOB models that can portray a dependency in past trade transactions are superior to models with the assumption of an infinitesimal tick size seen in the arithmetic Brownian motion model, as can be seen from the empirical results in He and Swishchuk, (2019), Swishchuk et al., (2019) and Swishchuk and Huffman, (2020). Many more studies have provided similar evidence, where some in-depth examples can be found in Cartea et al., 2018b , Cartea et al., 2018a , Mäkinen et al., (2019), Sjogren and DeLise, (2021), Gašperov and Kostanjčar, (2022) and many more.
As well as improving the price process modeling of LOBs, recent major advancements in Reinforcement Learning (RL), particularly in Deep RL, have enabled the formulation of a more state-of-the-art framework for solving problems in algorithmic and HFT. The most commonly used framework for studying these types of problems has normally been the AS model from Avellaneda and Stoikov, (2008), with numerous examples for various different types of trading problems such as liquidation, acquisition, MM, pairs-trading and statistical arbitrage given in Cartea et al., (2015) under the SOC setting. These methods, however, are not very robust and an alternative approach that has recently increased in popularity would be to find near-optimal controls under a deep RL framework, which is essentially an approximate solution deviating from the optimal solution by a small amount.
@@ 2 Deep Reinforcement Learning
SOC has often been the standard framework to solve algorithmic and HFT trading problems, with some important works including Bertsimas and Lo, (1998) in dealing with execution costs, Bouchard et al., (2011) where a general impulse control approach is applied to algorithmic trading problems, Fodra and Pham, (2015) applies optimal control to a HFT problem under a Markov Renewal approach, Cartea et al., (2015) gives a whole range of algorithmic and HFT problems such as acquisition, liquidation, MM, volume imbalance, statistical arbritrage, pairs trading, and in Guéant, (2017) a thorough theoretical overview is given for the optimal MM problem. In more recent times, the focus has shifted from SOC to deep RL for solving problems in algorithmic trading and HFT. One can find some recent general trading examples with applications in Li et al., (2019), Zhang et al., (2019), Théate and Ernst, (2021), and Cartea et al., (2021). This shift is mainly due to the inherent advantages in the RL space and we will now discuss three main reasons related to the kind of problems seen in algorithmic and HFT:
@@
    Model Uncertainty: SOC generally requires a well-defined model of the market dynamics, transition probabilities and reward structure. However, in reality, modeling financial markets is extremely complex due to the presence of many latent variables which could include sentiment, order flow information, liquidity, market microstructure noise, trading interactions and many more. RL does not require a well-defined model of these dynamics, and it can learn optimal strategies from interacting directly with the market. Thus, RL learns to continuously adapt to the markets highly unpredictable stochastic nature, which is essential for developing dynamic systems requiring flexible models.
@@ 4 Results
In order to perform our analysis we first trained our model on our simulated data, which was constructed based on the analysis in section 3. Please see below, in table 1, the parameter values used. The majority of these parameters were purely picked for illustrative purposes to assess how the deep RL algorithm could learn and perform, although some strategic selections were made. We kept the time step small as we would like to assess the strategy from a HFT standpoint. We set the maximum inventory parameter $q=5$ to avoid large positions significantly influencing the results. We will show results where we picked $20\%$ for our non-adverse fill probability, which is in line with empirical results in Lalor and Swishchuk, 2024b for estimating this probability based on some of the most liquid futures contracts listed on the Chicago Mercantile Exchange. We will then also proceed to show two sets of results, one with adverse fills and one without, in order to portray how significant the effect adverse fills can be on the overall results. We set the spread between the bid and the ask to $\Delta=0.01$ which is quite common throughout the majority of the day in many liquid financial products and we set the running inventory penalty to $0.001$, which had quite a bit of emphasis on the test results as inventory is generally very low. For our in-sample results, we set the number of simulations to $10^{6}$ for the training set, which lead to us training 1000 episodes. For our out-of-sample results, we then tested our training results on 200 unseen episodes. Computationally, we performed our analysis on python and for the SAC analysis part we made use of the Stable Baselines3 python package, developed by Raffin et al., (2020), which was also used by Gašperov and Kostanjčar, (2022) in their optimal MM problem.
@@
To summarize the training results, we provide visuals in figures 3 and 4, where the left and right figures portray these results by excluding (left) and including (right) adverse fills. It’s quite obvious that including this stylized fact, which is often overlooked in the literature, significantly effects the training results. In the training phase, the agent interacts with the trading environment by taking actions based on the policy. Following every action the MM agent takes, a reward is received after which they transition to the next state. The agents policy is then updated based on its reward function and its state transitions through the SAC algorithm, which we described in section 3. In figure 3, we show some results over the first training episode to portray how the optimal MM strategy evolved there. In the top subplots, we show a snapshot of the simulated bid and ask prices, along with the trade executions, then we show the simulated inventories in the second subplot row, and in the last subplot row we show the simulated cumulative rewards. Unsurprisingly, this agent has a negative cumulative reward at the beginning of the training process as it has not been able to figure out the optimal actions to take yet. As we get through more episodes, this slowly improves. To portray this improvement, we show the cumulative results over all 1000 episodes, see the histogram in figure 4, which shows the frequency of every terminal reward over all these episodes. It is quite apparent that the results are mostly positive when adverse fills are excluded, with some heavy left tail results with extreme negative rewards. When adverse fills are included, these results are still quite negative. Thus, although the agent has still improved in the training phase, adverse fills can still negatively hamper the results. This will become more evident when we move to the testing stage next.
@@
Next, we discuss the results from our test sample, $X_{test}$, which was conducted over 200 episodes. This step is crucial as it helps us determine whether our agent has learned to perform in this type of trading environment and whether a similar performance can be achieved on out-of-sample data. During the testing phase, the agent’s policy is fixed, which means it does not update its policy based on any rewards received. The environment we used under our training and testing sets are the same, which is based on our analysis in section 3, and this ensures unbiased results in machine learning terms. However, to assess the real-world deployment capabilities of any optimal MM problem, it would be prudent to also test the results of the training model in different types of market regimes, in particular regimes that were unseen in the training data, as this could significantly alter the out-of-sample results. Markets regularly change from one regime to another, which can often be measured through metrics such as the implied volatility. One should also test the robustness of any optimal MM strategy like this one, which can be done by changing the parameter values given in table 1.
@@
To visualize the testing results, we now provide similar visuals as we just did above for the training results, in figures 5 and 6. It is quiet obvious that the results have significantly improved in comparison to their respective training counterparts (in terms of the exclusion/inclusion of adverse fills), in particular regarding the left tail events present in the training phase, and the histogram shows that the cumulative rewards obtained are much more positive or less negative overall in each episode. In figure 5, we again show a snapshot of the simulated bid and ask prices in the top subplot row, along with the trade executions, then we show the simulated inventory in the second subplot row, and in the last subplot row we show the simulated cumulative reward. One thing that is very apparent is that the deep RL algorithm has learned it is best to not let the inventory deviate from -1 or 1. Thus, keeping a small position is optimal in this scenario. This likely helped to reduce the left tail events as the variables such as adverse fills have less of an effect when they occur if the inventory is low, while this outcome was also likely influenced by the inventory penalty term, $\alpha$. In figure 6, we can see that the cumulative terminal reward after each episode is much improved, thus under this specific framework the algorithm has learned to consistently generate positive or less negative rewards.
@@
Figure 3: Training results from the first training episode, excluding (left) and including (right) adverse fills.
@@
Figure 4: Histogram of cumulative rewards over all 1000 training episodes, excluding (left) and including (right) adverse fills.
@@
Figure 5: Testing results from the last testing episode, excluding (left) and including (right) adverse fills.
@@
Figure 6: Histogram of cumulative rewards over all 200 testing episodes, excluding (left) and including (right) adverse fills.
@@
As good as the above results may seem (in terms of learning to generate more positive or less negative cumulative rewards), we would like to stress some of their limitations, as much recent work has shed more light on the limitations of trading strategy back-test results under diffusion pricing models, in particular in a setting where many limit orders are placed, as in this MM strategy. We will now list some of the general limitations, where a great overview was given in Law and Viens, (2019). We did try to overcome some of these limitations by making adjustments to our optimal MM strategy, which we will also highlight again. These limitations and adjustments are as follows:
@@
  • In these types of frameworks, price and order arrivals are assumed to be independent. In order to tackle this problem, we made sure to include adverse fills in the simulation process as in Lalor and Swishchuk, 2024b . Excluding adverse fills has the ability to lead to “large phantom gains”, as described in Law and Viens, (2019), as MMs as liquidity providers are often on the wrong side of the trade. The empirical results in section 1.1 in Lalor and Swishchuk, 2024b also shows this, as well through the results we just showed in figures 3-6 above. Thus, including adverse fills has the ability to bring the results a lot closer to reality but still with some limitations, as it doesn’t exactly mimic how trading in live markets actually occurs, which is purely discrete based on specific LOB events. An overview of these events, under a reduced-form LOB model, can be seen in table 2 in Law and Viens, (2019), which was also applied to the discrete non-diffusive optimal MM deep RL problem in Gašperov and Kostanjčar, (2022).
@@ 5 Conclusions and Future Recommendations
In terms of recommendations for future research, we would like to highlight that if one could employ a full LOB model into the trading environment, as briefly described in Law and Viens, (2019), this would be a big step forward in improving the applicability of these types of models. In Gašperov and Kostanjčar, (2022) an attempt is made to create a reduced-form LOB model, where they also cite Law and Viens, (2019), which deploys a discrete-time multivariate linear Hawkes process model based on eight of the twelve events in the reduced-form LOB system. Other, and probably more simple model recommendations that we would recommend includes developing models where the spread is not fixed, computing stochastic non-adverse fill probabilities and stochastic volatility coefficients related to all the jump-diffusion coefficients, which may be related to either the semi-Markov or Hawkes models. In terms of the different types of trading problems out there, one could also try to apply this deep RL setting to the trading problems related to Liquidation, Acquisition, Statistical Arbitrage, Volume Imbalance and Pairs Trading as examples, which are some of the heavily studied trading problems under the standard SOC framework in Cartea et al., (2015).
Acknowledgments
We would like to thank MITACS and NSERC for research funding.